OpenStudy (anonymous):

273 g = _________________________ mg

OpenStudy (akashdeepdeb):

1g = 1000mg 273g = 273 * 1000mg = 273000mg !! ^_^ :D

OpenStudy (anonymous):

multiply * 1000

OpenStudy (anonymous):

= 273000

OpenStudy (anonymous):

thanks

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

273*1000 =273000mg